OpenStudy (anonymous):

−5(1 − 5x) + 5(−8x − 2) = −4x − 8x

OpenStudy (cj49):

-5+25x-40x-10=-4x-8x

OpenStudy (cj49):

-5+37x-40x-10=0 -3x=15 x=-15/3 x=-5

OpenStudy (anonymous):

Thanks :)

OpenStudy (cj49):

My Pleasure c: